17. Unpolarized light passes through a system of two polarizers. What would the angle between the polarizers have to be in order to have 15% of the original intensity of light pass through the system?
Answer #1

The amount of light passing through a system of two polarizers can be quantitatively described by applying Malu's cosine squared law.

I = I0*\cos2.\theta...........(1)

Where , I =intensity of light passing through the analyzer (total amount of light),

I0 = the intensity of light that is incident upon the polarizer

  \theta= angle between the transmission axes of the polarizer and analyzer

from question , 0.15 I0 = I

put the value of equation ....(1)

COS\theta= \sqrt{} (I / I0)= \large \sqrt{} 0.15=0.387

COs(0.387)=\large \theta

\large \theta=67.21 degree

\large \therefore the angle between the polarizers is 67.21 degree